# Rheology of dense vibrated granular flows: non-monotonic response controlled by granular temperature

### v1

@@ Abstract
We study the rheology of dense granular materials subjected to vertical vibration using numerical simulations of a stress-imposed vane rheometer. The effective viscosity increases with confining pressure, decreases with vibration amplitude, and exhibits a non-monotonic dependence on frequency: weakening is observed at intermediate frequencies but is lost at high frequencies. We show that the rheological response is governed by the balance between grain-scale agitation energy and the stabilizing effect of confinement. This framework reconciles previously observed trends in viscosity and friction weakening and emphasizes the central role of energy injection and dissipation in determining granular flow properties under vibration.
@@ I Introduction
Granular materials are ubiquitous in natural environment and applications [andreotti2013granular, puglisi2014transport]. Depending on the external forcing, they can exhibit both solid-like or fluid-like behaviour [jaeger1996granular, eshuis2007phase]. The rheology of granular materials in shear cells or flowing over inclined planes has been extensively investigated [bagnold1954experiments, jop2006constitutive, forterre2008flows, da2005rheophysics, bouzid2013nonlocal, fall2015dry, deboeuf2023cohesion]. At the same time, the role of vibration in determining granular flow properties has been highlighted in several theoretical, numerical and experimental studies [melosh1996dynamical, dijksman2011jamming, wortel2014rheology, giacco2015dynamic, gnoli2016unified, gnoli2018controlled, leopoldes2020triggering, Plati2021Getting, Clark2023, irmer2025granular, d2025rheological]. External vibrations usually increase the flowability of a granular system and this can have dramatic consequences as in the case of earthquake-triggered landslides [garcia2010assessment, leopoldes2020triggering] or can be exploited to facilitate the manipulation of granular systems in applications [wang2021investigation]. In laboratory experiments, vertical vibration is known to reduce to zero the yield stress and produce non-monotonic flow curves [dijksman2011jamming]. It may also induce a transition from thinning to collisional thickening [gnoli2016unified].
@@
The general setting for experiments or simulations to probe the rheology of granular materials under sinusoidal vibration considers the role of three main parameters: the vibration amplitude $A$, the vibration frequency $f$ and the confining pressure $p$. Depending on the specific setup, different quantities can be used as proxies of the “resistance to flow” of the granular system. When the shear rate $\dot{\gamma}$ is imposed, one usually measures the macroscopic friction coefficient $\mu=\tau/p$ where $\tau$ is the resulting shear stress. Conversely, if the shear stress is imposed, the flow is characterized by the effective viscosity $\eta=\tau/\dot{\gamma}$. Non-vibrated granular systems have been extensively studied by focusing on the so-called flow curves $\mu(\dot{\gamma})$ and $\eta(\tau)$ [fall2015dry, forterre2008flows, DaCruz2002], however - when vibrations are taken into account - the goal is to characterize the flowability of the material either through $\mu(\dot{\gamma},A,f,p)$ or $\eta(\tau,A,f,p)$. To restrict the parameter space, it is sometime useful to consider $\dot{\gamma}$ and $\tau$ fixed, focusing on $\mu(A,f,p)$ and $\eta(A,f,p)$. In recent years, many experimental and numerical observations on these quantities have been reported [gnoli2018controlled, Plati2021Getting, Clark2023, irmer2025granular].
@@
In Ref. [gnoli2018controlled], the effective viscosity of a dense, weakly fluidized granular system was investigated experimentally and numerically by varying the driving amplitude $A$ and frequency $f$ under a fixed confining load. It was found that $\eta$ decreases monotonically with $A$, whereas its dependence on $f$ is non-monotonic: the viscosity decreases until a minimum at a frequency $f^{*}$ that shifts to lower values as the dissipative properties of the medium increase. A subsequent work [Plati2021Getting] showed that these trends are linked to the behavior of the average kinetic energy of the system (i.e. the granular temperature, see Sec. IIA for its definition), which exhibits a maximum at $f^{*}$. In Fig. 1, we show the granular temperature as a function of $f$ and $A$ obtained extending the numerical dataset of Ref. [Plati2021Getting]. It increases monotonically with $A$ when the frequency is fixed and exhibits a non-monotonic trend with $f$ when the amplitude is fixed, which correlates with the behavior of the effective viscosity obtained in experiments [gnoli2018controlled]. These results suggest that rheology is primarily controlled by the efficiency with which the medium absorbs energy from external vibration. As discussed in [Plati2021Getting], while the non-monotonic dependence on $f$ resembles a resonance phenomenon, the fact that $f^{*}$ depends on dissipation rules out a standard resonance, in which the peak position is fixed by the elastic properties of the system. The proposed interpretation instead attributes the non-monotonicity to the competition between the energy injected by the vibration mechanism, and the dissipation rate of the medium (which is related to the grain-grain collision rate). Indeed, both these competing effects are enhanced by the driving frequency, producing a non-monotonic trend.
@@
Figure 1: Granular temperature as a function of the oscillation amplitude $A$ and frequency $f$. The granular temperature increases monotonically as a function of AA at any fixed frequency. Meanwhile, for any fixed value of $A$, it exhibits a non-monotonic behavior, reaching a maximum at an intermediate frequency. This plot is obtained by extending the numerical results of Ref. [Plati2021Getting] to a broader range of $A$ and $f$ values.
@@
More recently, Clark et al. [Clark2023] examined the friction coefficient $\mu(A,f,p)$ and observed a similar non-monotonic trend with frequency. However, rather than focusing on the frequency dependence, they analysed the behavior of the minimum friction coefficient $\mu_{\min}(A,p)$ which was found to be controlled by the variable $A^{2}/p$. This scaling was motivated by the idea that friction weakening requires vibration amplitudes large enough to disrupt the underlying force network: while shaking at sufficiently high acceleration is enough to break individual contacts, an additional threshold in $A^{2}/p$ is necessary to destabilize mesoscale structures such as force chains.
@@
In our study we go beyond this interpretation, based on the evidence that, regardless of whether the thresholds in shaking acceleration and $A^{2}/p$ are exceeded, friction weakening is always lost at sufficiently high $f$. Following the interpretation of [Plati2021Getting], we highlight that the driving frequency governs the efficiency of energy transfer into the granular medium, which in turn controls the system’s rheological properties. The centrality of energy transfer has also been evidenced in more recent studies, focusing on local rheological responses [irmer2025granular]. We therefore emphasize that any comprehensive description of friction weakening must include a detailed analysis of a wider frequency range.
@@
In this article, we build on these findings to provide a unified framework for understanding the interplay between dissipation, energy injection, and rheological response in realistic vibrated granular systems. In Sec. II we describe our numerical setup which is a realistic in silico reproduction of the experimental system for vane rheometry studied in [gnoli2018controlled, Plati2021Getting]. In the same section we also provide an overview of the rheological response for different shaking parameters and confining pressures. This is done by discussing the behavior of the effective viscosity $\eta(A,f,p)$. This analysis is complemented by the introduction of a minimal model (detailed in Appendix C), which accounts for the combined effect of shaking amplitude and confining pressure. In Sec. III, we show that the key variable controlling $\eta$ involves the ratio between the average kinetic energy of the grains and the confining pressure. There we also discuss a physical interpretation of these results based on the competition between two energy scales, the first related to grain motion and second related to the stability of local structures. In Sec. IV, we conclude our paper by discussing the link between our results and previous studies; we also provide a brief outlook for future developments of this work.
@@ II Effective viscosity of the granular medium
Our numerical setup is the in silico reconstruction of the experimental apparatus used in [gnoli2018controlled, Plati2021Getting] and it has been shown to provide results in agreement with experiments in previous studies [plati2019dynamical, Plati2021Getting]. We use the Discrete Element Method (DEM) [Cundall79] to perform numerical simulations of a monodispersed granular system with $N=2600$ particles of radius $r=2$ mm and mass $m=0.27$ g confined in a cylinder with a conical-shaped base and covered by a circular lid of mass $M$ and radius $R=45$ mm, which matches the area of the cylinder’s round face. The whole system feels gravity acceleration $g$. A rectangular-shaped vane, able to rotate around a fixed vertical axis, is immersed in the granular medium and subjected to a constant torque $\mathcal{T}=0.006$ Nm. The granular material is subjected to a sinusoidal vertical displacement $\Delta z(t)=A\sin(2\pi ft)$ which is imposed on the system’s container. The torque value is chosen so that the vane does not move in the absence of vibration. We implement our numerical simulations through the LAMMPS package [Plimpton1995, Plimpton2022, LammpsSiteGran] which uses the Hertz-Mindlin contact model [Zhang2005, LammpsSiteGran] to describe grain-grain, grain-vane, grain-lid and grain-wall interactions. More technical details about our numerical methods are provided in Appendix A.
Figure 2: a) Effective viscosity as a function of the rescaled acceleration (varied through $f$) for three shaking amplitudes. Experimental data from [gnoli2018controlled]. b) Analogous results by numerical simulations. Each point is obtained by averaging over 3 statistically independent runs. c) Kinetic energy of the granular medium as a function of $\Gamma$ (varied through $f$) for three shaking amplitudes (simulations performed without the rotating vane). d) Effective viscosity as a function of the granular temperature in numerical simulations. The decreasing trend is well approximated by the scaling $\eta\propto K^{\alpha}$ with $\alpha=-2$.
@@ II.1 Non-monotonic behavior as a function of the shaking frequency
Following a microrheology approach, we measure the angular velocity $\Omega(t)$ of the vane as a function of time and explore its behaviour varying the frequency $f$ and the amplitude $A$ of the imposed shaking. In view of this, we introduce the rescaled acceleration $\Gamma=(2\pi f)^{2}A/g$.
@@
We have checked that, for all the explored combinations of $A$ and $f$, our system always lies in a weakly fluidized regime which means that grains move inside the cage formed by the surrounding particles without experiencing substantial rearrangement except for that induced by the vane’s motion. Once the vibration and the torque are switched on, after a $t_{0}$-long initial transient, the vane reaches a stationary state with an average velocity $\langle\Omega\rangle=1/T\int_{t_{0}}^{t_{0}+T}\Omega(t)dt$ which can be thought as a proxy of the shear rate. For all our simulations we have $t_{0}=13$ s and $T=33$ s. We define the effective viscosity of the granular medium as $\eta=\mathcal{T}/|\langle\Omega\rangle|$. The choice of this variable is motivated by our previous experiments [gnoli2018controlled] showing a non-monotonic behaviour of $\langle\Omega\rangle$ and in turn of $\eta$ as a function of the frequency. In Fig. 2a-b we provide a comparison between experimental and numerical data exhibiting the resulting non-monotonic dependence of $\eta$ as a function of $\Gamma$ (varied through $f$) for different $A$. As highlighted by the vertical lines in the plots, the decreasing branch of the $\eta(\Gamma)$ curves can be explained in terms of the fluidization threshold around $\Gamma=1$ [gnoli2018controlled], where the shaking acceleration becomes equal to $g$. The physical mechanism underlying the viscosity strengthening at larger $\Gamma$ has been discussed in a previous work [Plati2021Getting], where we showed that it is linked to the behavior of the average kinetic energy of the granular medium in the steady state $K=1/T\int_{t_{0}}^{t_{0}+T}K_{\text{inst}}(t)dt$, where $K_{\text{inst}}(t)=\frac{m}{2N}\sum_{i}v_{i}^{2}(t)$ and $v_{i}(t)$ is the instantaneous speed of the grain $i$. The variable $K$ is usually called granular temperature (or directly proportional to it through the number of degrees of freedom which is constant here) and displays itself a non-monotonic dependence on the vibration frequency as shown in Fig. 2c. This phenomenon originates from the competing effects of injected vibration energy and (inelastic) grain-grain collision rate which both increase with $f$ [Plati2021Getting]. A more detailed numerical analysis supporting this interpretation based on energy dissipation can be found in Appendix B. The behavior of $\eta$ as a function of the driving parameters $A$ and $f$ is then effectively controlled by $K$, as highlighted in Fig. 2d. We point out here that to investigate the effect of the granular temperature on $\eta$ without considering trivial correlations due to the blade motion, we measured $K$ in simulations without the rotating blade.
@@ II.2 Effect of normal pressure
We now focus on the behaviour of $\eta$ as a function of $\Gamma$ (varied through $f$) for different couples of $\{A,p\}$ where $p=Mg/(\pi R^{2})$ is the normal pressure due to the lid’s weight. In Fig. 3a, we observe that all curves decrease up to the fluidization threshold and then reach a minimum $\eta_{\text{min}}$ (highlighted with a larger mark) at $\Gamma=\Gamma_{\text{m}}$ before showing a gradual increase. As expected, increasing $A$ reduces the effective viscosity while raising $p$ has the opposite effect. In order to first focus on the most fluidized regime, we get rid for the moment of the $\Gamma$ dependence considering the behavior of $\eta_{\text{min}}(A,p)$. In Fig. 3b, we show that $\eta_{\text{min}}(A,p)$ is mainly controlled by the variable $A^{2}/p$. Remarkably, a similar data collapse has been previously observed also for the macroscopic friction coefficient in simulations of 2D sheared granular materials under vertical vibrations [Clark2023]. Our results generalize this observation to the effective viscosity of a 3D granular system probed through stress-imposed microrheology. The role of the variable $A^{2}/p$ was explained in Ref. [Clark2023] based on physical mechanisms involving the rupture of force chains in the bulk of the granular medium. In Appendix C we discuss a simple mechanical model which is able to reproduce the $A^{2}/p$ scaling and allows us to trace it back to the interplay between energy injection and non-linear elastic deformation. This shows that force chain breaking is not necessary to explain the observed scaling.

### v2

# arXiv:2511.14859v2[cond-mat.soft] 02 Mar 2026

@@ Abstract
We study the rheology of dense granular materials subjected to vertical vibration by using numerical simulations of a stress-imposed vane rheometer. The effective viscosity increases with confining pressure, decreases with vibration amplitude, and exhibits a non-monotonic dependence on frequency: weakening is observed at intermediate frequencies but is lost at high frequencies. We show that the rheological response is governed by the balance between grain-scale agitation energy and the stabilizing effect of confinement. This framework reconciles previously observed trends in viscosity and friction weakening and emphasizes the central role of energy injection and dissipation in determining granular flow properties under vibration.
@@ I Introduction
Granular materials are ubiquitous in natural environment and applications [3, 31]. Depending on the external forcing, they can exhibit both solid-like or fluid-like behaviour [24, 15]. The rheology of granular materials in shear cells or flowing over inclined planes has been extensively investigated [4, 25, 17, 11, 6, 16, 13]. At the same time, the role of vibration in determining granular flow properties has been highlighted in several theoretical, numerical and experimental studies [27, 14, 22, 34, 19, 21, 20, 26, 29, 7, 23, 9]. External vibrations usually increase the flowability of a granular system and this can have dramatic consequences as in the case of earthquake-triggered landslides [18, 26] or can be exploited to facilitate the manipulation of granular systems in applications [33]. In laboratory experiments, vertical vibration is known to reduce to zero the yield stress and produce non-monotonic flow curves [14]. It may also induce a transition from thinning to collisional thickening [21].
@@
The general setting for experiments or simulations to probe the rheology of granular materials under sinusoidal vibration considers the role of three main parameters: the vibration amplitude $A$, the vibration frequency $f$ and the confining pressure $p$. Depending on the specific setup, different quantities can be used as proxies of the “resistance to flow” of the granular system. When the shear rate $\dot{\gamma}$ is imposed, one usually measures the macroscopic friction coefficient $\mu=\tau/p$ where $\tau$ is the resulting shear stress. Conversely, if the shear stress is imposed, the flow is characterized by the effective viscosity $\eta=\tau/\dot{\gamma}$. Non-vibrated granular systems have been extensively studied by focusing on the so-called flow curves $\mu(\dot{\gamma})$ and $\eta(\tau)$ [16, 17, 10], however - when vibrations are taken into account - the goal is to characterize the flowability of the material either through $\mu(\dot{\gamma},A,f,p)$ or $\eta(\tau,A,f,p)$. To restrict the parameter space, it is sometime useful to consider $\dot{\gamma}$ and $\tau$ fixed, focusing on $\mu(A,f,p)$ and $\eta(A,f,p)$. In recent years, many experimental and numerical observations on these quantities have been reported [20, 29, 7, 23].
@@
In Ref. [20], the effective viscosity of a dense, weakly fluidized granular system was investigated experimentally and numerically by varying the driving amplitude $A$ and frequency $f$ under a fixed confining load. It was found that $\eta$ decreases monotonically with $A$, whereas its dependence on $f$ is non-monotonic: the viscosity decreases until a minimum at a frequency $f^{*}$ that shifts to lower values as the dissipative properties of the medium increase. A subsequent work [29] showed that these trends are linked to the behavior of the average kinetic energy of the system (i.e. the granular temperature, see Sec. IIA for its definition), which exhibits a maximum at $f^{*}$. In Fig. 1, we show the granular temperature as a function of $f$ and $A$ obtained by extending the numerical dataset of Ref. [29]. It increases monotonically with $A$ when the frequency is fixed and exhibits a non-monotonic trend with $f$ when the amplitude is fixed, which correlates with the behavior of the effective viscosity obtained in experiments [20]. These results suggest that rheology is primarily controlled by the efficiency with which the medium absorbs energy from external vibration. As discussed in [29], while the non-monotonic dependence on $f$ resembles a resonance phenomenon, the fact that $f^{*}$ depends on dissipation rules out a standard resonance, in which the peak position is fixed by the elastic properties of the system. The proposed interpretation instead attributes the non-monotonicity to the competition between the energy injected by the vibration mechanism, and the dissipation rate of the medium (which is related to the grain-grain collision rate). Indeed, both these competing effects are enhanced by the driving frequency, producing a non-monotonic trend.
@@
Figure 1: Granular temperature as a function of the oscillation amplitude $A$ and frequency $f$. The granular temperature increases monotonically as a function of AA at any fixed frequency. Meanwhile, for any fixed value of $A$, it exhibits a non-monotonic behavior, reaching a maximum at an intermediate frequency. This plot is obtained by extending the numerical results of Ref. [29] to a broader range of $A$ and $f$ values.
@@
More recently, Clark et al. [7] examined the friction coefficient $\mu(A,f,p)$ and observed a similar non-monotonic trend with frequency. However, rather than focusing on the frequency dependence, they analysed the behavior of the minimum friction coefficient $\mu_{\min}(A,p)$ which was found to be controlled by the variable $A^{2}/p$. This scaling was motivated by the idea that friction weakening requires vibration amplitudes large enough to disrupt the underlying force network: while shaking at sufficiently high acceleration is enough to break individual contacts, an additional threshold in $A^{2}/p$ is necessary to destabilize mesoscale structures such as force chains.
@@
In our study we go beyond this interpretation, based on the evidence that, regardless of whether the thresholds in shaking acceleration and $A^{2}/p$ are exceeded, friction weakening is always lost at sufficiently high $f$. Following the interpretation of [29], we highlight that the driving frequency governs the efficiency of energy transfer into the granular medium, which in turn controls the system’s rheological properties. The centrality of energy transfer has also been evidenced in more recent studies, focusing on local rheological responses [23]. We therefore emphasize that any comprehensive description of friction weakening must include a detailed analysis of a wider frequency range.
@@
In this article, we build on these findings to provide a unified framework for understanding the interplay between dissipation, energy injection, and rheological response in realistic vibrated granular systems. In Sec. II we describe our numerical setup which is a realistic in silico reproduction of the experimental system for vane rheometry studied in [20, 29]. In the same section we also provide an overview of the rheological response for different shaking parameters and confining pressures. This is done by discussing the behavior of the effective viscosity $\eta(A,f,p)$. This analysis is complemented by the introduction of a minimal model (detailed in Appendix C), which accounts for the combined effect of shaking amplitude and confining pressure. In Sec. III, we show that the key variable controlling $\eta$ involves the ratio between the average kinetic energy of the grains and the confining pressure. There we also discuss a physical interpretation of these results based on the competition between two energy scales, the first related to grain motion and second related to the stability of local structures. In Sec. IV, we conclude our paper by discussing the link between our results and previous studies; we also provide a brief outlook for future developments of this work.
Figure 2: Sketch of the experimental setup used in Ref. [20, 29]. A rotating vane with an imposed constant torque is immersed in a granular medium confined in a conical-shaped container covered by a circular lid. Our numerical setup consists of the in silico reconstruction of this apparatus (details in the text).
Figure 3: a) Effective viscosity as a function of the rescaled acceleration (varied through $f$) for three shaking amplitudes. Experimental data from [20]. b) Analogous results by numerical simulations. Each point is obtained by averaging over 3 statistically independent runs. c) Kinetic energy of the granular medium as a function of $\Gamma$ (varied through $f$) for three shaking amplitudes (simulations performed without the rotating vane). d) Effective viscosity as a function of the granular temperature in numerical simulations. The decreasing trend is well approximated by the scaling $\eta\propto K^{\alpha}$ with $\alpha\in[-2,-1.5]$. Effective viscosities are obtained from angular velocities expressed in degrees per seconds.
@@ II Effective viscosity of the granular medium
Our numerical setup is the in silico reconstruction of the experimental apparatus used in [20, 29] and it has been shown to provide results in agreement with experiments in previous studies [28, 29] (see Fig. 2). We use the Discrete Element Method (DEM) [8] to perform numerical simulations of a monodispersed granular system with $N=2600$ particles of radius $r=2$ mm and mass $m=0.27$ g confined in a cylinder with a conical-shaped base and covered by a circular lid of mass $M$ and radius $R=45$ mm, which matches the area of the cylinder’s round face. The whole system feels gravity acceleration $g$. A rectangular-shaped vane, able to rotate around a fixed vertical axis, is immersed in the granular medium and subjected to a constant torque $\mathcal{T}=0.006$ Nm. The granular material is subjected to a sinusoidal vertical displacement $\Delta z(t)=A\sin(2\pi ft)$ which is imposed on the system’s container. The torque value is chosen so that the vane does not move in the absence of vibration. We implement our numerical simulations through the LAMMPS package [30, 32, 1] which uses the Hertz-Mindlin contact model [35, 1] to describe grain-grain, grain-vane, grain-lid and grain-wall interactions. More technical details about our numerical methods are provided in Appendix A.
@@ II.1 Non-monotonic behavior as a function of the shaking frequency
By following a microrheology approach, we measure the angular velocity $\Omega(t)$ of the vane (always expressed in degrees per seconds) as a function of time and explore its behaviour by varying the frequency $f$ and the amplitude $A$ of the imposed shaking. In view of this, we introduce the rescaled acceleration $\Gamma=(2\pi f)^{2}A/g$.
@@
We have checked that, for all the explored combinations of $A$ and $f$, our system always lies in a weakly fluidized regime which means that grains move inside the cage formed by the surrounding particles without experiencing substantial rearrangement except for that induced by the vane’s motion. Once the vibration and the torque are switched on, after a $t_{0}$-long initial transient, the vane reaches a stationary state with an average velocity $\langle\Omega\rangle=1/T\int_{t_{0}}^{t_{0}+T}\Omega(t)dt$ which can be thought as a proxy of the shear rate. For all our simulations we have $t_{0}=13$ s and $T=33$ s. We define the effective viscosity of the granular medium as $\eta=\mathcal{T}/|\langle\Omega\rangle|$. The choice of this variable is motivated by our previous experiments [20] showing a non-monotonic behaviour of $\langle\Omega\rangle$ and in turn of $\eta$ as a function of the frequency. In Fig. 3a-b we provide a comparison between experimental and numerical data exhibiting the resulting non-monotonic dependence of $\eta$ as a function of $\Gamma$ (varied through $f$) for different $A$. As highlighted by the vertical lines in the plots, the decreasing branch of the $\eta(\Gamma)$ curves can be explained in terms of the fluidization threshold around $\Gamma=1$ [20], where the shaking acceleration becomes equal to $g$. The physical mechanism underlying the viscosity strengthening at larger $\Gamma$ has been discussed in a previous work [29], where we showed that it is linked to the behavior of the average kinetic energy of the granular medium in the steady state $K=1/T\int_{t_{0}}^{t_{0}+T}K_{\text{inst}}(t)dt$, where $K_{\text{inst}}(t)=\frac{m}{2N}\sum_{i}v_{i}^{2}(t)$ and $v_{i}(t)$ is the instantaneous speed of the grain $i$. The variable $K$ is usually called granular temperature (or directly proportional to it through the number of degrees of freedom which is constant here) and displays itself a non-monotonic dependence on the vibration frequency as shown in Fig. 3c. This phenomenon originates from the competing effects of injected vibration energy and (inelastic) grain-grain collision rate which both increase with $f$ [29]. A more detailed numerical analysis supporting this interpretation based on energy dissipation can be found in Appendix B. The behavior of $\eta$ as a function of the driving parameters $A$ and $f$ is then effectively controlled by $K$, as highlighted in Fig. 3d. The decreasing trend is well approximated by the power-law scaling $\eta\propto K^{\alpha}$, with $\alpha\in[-2,-1.5]$ (see the guides for the eye). We finally point out here that to investigate the effect of the granular temperature on $\eta$ without considering trivial correlations due to the blade motion, we measured $K$ in simulations without the rotating blade.
@@ II.2 Effect of normal pressure
We now focus on the behaviour of $\eta$ as a function of $\Gamma$ (varied through $f$) for different couples of $\{A,p\}$ where $p=Mg/(\pi R^{2})$ is the normal pressure due to the lid’s weight. In Fig. 4a, we observe that all curves decrease up to the fluidization threshold and then reach a minimum $\eta_{\text{min}}$ (highlighted with a larger mark) at $\Gamma=\Gamma_{\text{m}}$ before showing a gradual increase. As expected, increasing $A$ reduces the effective viscosity while raising $p$ has the opposite effect. In order to first focus on the most fluidized regime, we get rid for the moment of the $\Gamma$ dependence and consider the behavior of $\eta_{\text{min}}(A,p)$. In Fig. 4b, we show that $\eta_{\text{min}}(A,p)$ is mainly controlled by the variable $A^{2}/p$. Remarkably, a similar data collapse has been previously observed also for the macroscopic friction coefficient in simulations of 2D sheared granular materials under vertical vibrations [7]. Our results generalize this observation to the effective viscosity of a 3D granular system probed through stress-imposed microrheology. The role of the variable $A^{2}/p$ was explained in Ref. [7] based on physical mechanisms involving the rupture of force chains in the bulk of the granular medium. In Appendix C we discuss a simple mechanical model which is able to reproduce the $A^{2}/p$ scaling and allows us to trace it back to the interplay between energy injection and non-linear elastic deformation. This shows that force chain breaking is not necessary to explain the observed scaling.